### v1

# Do LLMs write like humans? Variation in grammatical and rhetorical styles

Abstract
Large language models (LLMs) are capable of writing grammatical text that follows instructions, answers questions, and solves problems. As they have advanced, it has become difficult to distinguish their output from human-written text. While past research has found some differences in surface features such as word choice and punctuation, and developed classifiers to detect LLM output, none has studied the rhetorical styles of LLMs.
Using several variants of Llama 3 and GPT-4o, we construct two parallel corpora of human- and LLM-written texts from common prompts. Using Douglas Biber’s set of lexical, grammatical, and rhetorical features, we identify systematic differences between LLMs and humans and between different LLMs. These differences persist when moving from smaller models to larger ones, and are larger for instruction-tuned models than base models. This demonstrates that despite their advanced abilities, LLMs struggle to match human styles, and hence more advanced linguistic features can detect patterns in their behavior not previously recognized.
1 Introduction
As large language models (LLMs) have advanced in recent years, from “stochastic parrots” to models evidently capable of performing complex tasks, most attention has focused on their reasoning performance: solving mathematical problems, writing code, evaluating arguments, diagnosing diseases, and so on (Wang et al., 2023; Huang and Chang, 2023). While past research has studied their mastery of basic grammar and vocabulary (Chang and Bergen, 2024), there is relatively little research on their language performance more generally: their ability to produce readable text in a variety of styles. Rather than exploring it in detail, commentators discuss the business and communication tasks that might be automated by LLMs with their writing ability, or consider the dangers of impersonation and misinformation facilitated by LLMs (Barman et al., 2024; Kumar et al., 2023; Kovács, 2024). As more and more writing tasks are automated, such problems appear inevitable.
However, the impression that LLMs write “like humans” is based primarily on qualitative evaluation of their output, not on thorough linguistic evaluation of their text. So far, quantitative comparisons have looked mainly at basic grammar and syntax (Chang and Bergen, 2024) or surface features such as word choice, punctuation, sentence length, and so on, finding evidence of some differences between human- and LLM-written text (Tang et al., 2024; Fröhling and Zubiaga, 2021; Muñoz-Ortiz et al., 2024; Herbold et al., 2023; Liang et al., 2024). Other work has used these features, or language models trained on sample texts, to classify LLM-written texts with varying degrees of success (Liu et al., 2024; Mosca et al., 2023; Wang et al., 2024). Though not definitive, these results suggest there are indeed structural differences between human- and LLM-written text.
In this study, we used several recent LLMs (OpenAI’s GPT-4o and GPT-4o Mini, and four variants of Meta Llama 3) to generate text from prompts drawn from a large, representative corpus of English, allowing us to directly compare the style of LLM writing to human writing. We find large differences in grammatical, lexical, and stylistic features, demonstrating that LLMs prefer specific grammatical structures and struggle to match the stylistic variation present in human writing. In Llama 3, where we are able to compare base models to instruction-tuned variants, we further see that the instruction tuning introduces more extreme grammatical differences, making them easier to distinguish from human writing and introducing features similar to those present in GPT-4o and GPT-4o Mini.
For example, the instruction-tuned LLMs used present participial clauses at 2 to 5 times the rate of human text, such as in this sentence from GPT-4o using two present participles: “Bryan, leaning on his agility, dances around the ring, evading Show’s heavy blows.” They also use nominalizations at 1.5 to 2 times the rate of humans, such as in this sentence from Llama 3 70B Instruct containing four: “These schemes can help to reduce deforestation, habitat destruction, and pollution, while also promoting sustainable consumption patterns.” On the other hand, GPT-4o uses the agentless passive voice at roughly half the rate as human texts—but in each case, the Llama base models use these features at rates more closely matching humans. This suggests that instruction tuning, rather than training the models to write even more like humans, instead trains them in a particular informationally dense, noun-heavy style, and limits their ability to mimic other writing styles.
These results demonstrate the value of attending to linguistic structure (morphosyntactic, functional, and rhetorical) in order to better understand the affordances and outputs of large language models. Since the rise of the Internet and the concurrent development of efficient processing architectures, language modeling has relied on relatively simple linguistic principles (i.e., sequences and context windows) as the availability of massive amounts of text allowed models trained on text to rapidly outperform older paradigms based on linguistic theory (Li, 2023); but linguistic theory can provide better ways to evaluate LLM output, just as improved benchmark problems can provide better ways to evaluate their reasoning ability. These results also demonstrate the limits of current LLMs in matching human language, showing that despite their apparent ability, they have measurable limitations compared to human authors.
2 Methods
2.1 Human-AI parallel corpora
For this study, we created two parallel corpora. The construction of each corpus was accomplished by:
  1. 1.
    assembling 12,000 human-authored texts in English representing a range of registers (or types);
  2. 2.
    selecting roughly the first 1,000 words of each text (where that span does not cross a section or chapter boundary);
  3. 3.
    splitting each string into approximately 500-word chunks (with the split-point being a sentence boundary as determined by a spaCy dependency parse);
  4. 4.
    passing the first chunk to six different LLMs and prompting the models to generate 500 more words of text in the same style, similar to the method used by Yang et al. (2024);
  5. 5.
    discarding all texts that contain a non-response or a nonsense response (like a string of punctuation) from any of the six LLMs.
The resulting corpora, then, contain the first chunk of human-authored text (the “germinal” text used for the LLM prompt) and seven variations of the second chunk: the original (human-authored) continuation of the first chunk, and continuations generated by six different LLMs. This workflow is illustrated in Figure 1. The figure also highlights the alignment of the second, parallel chunks, which enable comparisons between the writing of humans and LLMs, as well as comparisons among the writing produced by the LLMs, themselves.
Chunk 1LLM prompt Chunk 2 Llama 3 8B Llama 3 8B Instruct Llama 3 70B Llama 3 70B Instruct GPT-4o GPT-4o Mini
Figure 1: The LLM text generation workflow. The human text was split into two chunks of roughly 500 words; the first chunk was used to prompt an LLM to create text that was compared to the second human chunk.
The first of our parallel corpora uses the Corpus of Contemporary American English (COCA) as the source of its human-authored text and is called the COCA AI Parallel (CAP) corpus. As noted above, we wanted data representative of a broad range of sources and registers. COCA is a monitor corpus with over 1 billion words and balanced among eight registers: spoken, fiction, magazines, newspapers, academic journals, blogs, other web pages, and TV/movie subtitles (Davies, 2008). For our dataset, we filtered for texts containing a minimum of approximately 1,250 words, then randomly sampled 1,500 documents from each register to create a corpus of 12,000 texts.
Though its breadth and representativeness make COCA valuable for this task, it does have a key limitation: to allow distribution of otherwise copyrighted material, COCA alters the text by removing 10 words every 200 words. [^1] See https://www.corpusdata.org/limitations.asp for details on the procedure. This produces gaps and incomplete sentences in each 500-word chunk that might affect sentence parsing and feature extraction. For this reason, we constructed a second corpus, which is meant to approximate COCA’s composition.
The resulting dataset, the Human AI Parallel English (HAP-E) corpus, is drawn from six (as opposed to COCA’s eight) text-types: [^2] The data can be accessed from https://huggingface.co/datasets/browndw/human-ai-parallel-corpus.
  • Academic: samples were drawn from a 40,000+ document corpus of open-access academic articles published by Elsevier.
  • News: samples were drawn from a 100,000+ document corpus of news articles published online by U.S.-based news organizations.
  • Fiction: samples were drawn from novels and short stories in the public domain and available on Project Gutenberg.
  • Spoken: samples were drawn from a corpus of 100,000 podcast transcriptions. (Note that COCA’s spoken text-type comes from transcriptions of unscripted news shows.)
  • Blogs: samples were drawn from a corpus of 681,288 posts from blogger.com.
  • Television and Movie Scripts: samples were drawn from two different corpora of scripts, some of which were converted using OCR.
The sampling and processing of the HAP-E corpus followed the same five steps described at the beginning of this section. The word counts of the HAP-E and CAP corpora can be found in the Appendix, Tables LABEL:tbl:totals and LABEL:tbl:cap_totals.
2.2 LLM text generation
We used multiple versions of GPT-4o (OpenAI, 2024a, b) and Meta Llama 3 (Meta, 2024) to generate the parallel LLM texts, shown in Table 1. We chose these LLMs because they are popular, available within a reasonable budget, and generally considered to perform very well. Meta Llama 3 is available in both base models, which provide text completion, and instruction-tuned variants that use additional human feedback to optimize the models to follow instructions and answer questions; recent GPT variants are only available in instruction-tuned form. Llama 3’s availability in both forms allows us to observe the effects of instruction tuning.
Model Version Context Instruct?
GPT-4o Mini 2024-07-18 128K Yes
GPT-4o 2024-08-06 128K Yes
Llama 3 8B 8K No
Llama 3 8B-Instruct 8K Yes
Llama 3 70B 8K No
Llama 3 70B-Instruct 8K Yes
Table 1: Large language models used to generate text completions.
Instruction-tuned LLMs were given the following prompt, followed by the first chunk of human text:
In the same style, tone, and diction of the following text, complete the next 500 words, generate exactly 500 words, and note that the text does not necessarily end after the generated words:
Base Llama 3 variants without instruction tuning were simply given the first chunk of text. In each case, the LLMs generated roughly 500 words of additional output, continuing where the first chunk finished. This could be compared directly to the second chunk of human text. GPT-4o text was generated using OpenAI’s Batch API using its default settings; Llama 3 text was generated on a local GPU cluster using the vLLM package for Python, version 0.5.5 (Kwon et al., 2023).
As some LLMs refused to complete some prompts, or returned short, unusable answers, we removed responses shorter than 100 words, producing 8,290 HAP-E texts and 9,615 COCA texts with valid responses from all LLMs.
We designed this procedure for creating parallel data because our aims are focused more on description than they are on detection. Compared to other methods that only prompt LLMs with a title or topic (e.g. Wang et al., 2024), this approach gives the LLMs roughly 500 words of context so they can adapt their output and its style to the human text.
2.3 Feature extraction
To extract meaningful features from our corpus for training our classifiers, we used Douglas Biber’s tagset of 66 linguistic categories, gleaned from the many works Biber has authored and co-authored (Biber, 1988, 1995; Biber and Conrad, 2009). Biber’s tagset includes indices of lexical complexity (e.g., mean word length), but many more are raw linguistic features ranging from the lexical (e.g., thinking verbs) to the grammatical (e.g., adjectives before a noun). For example, features include:
  • Nominalizations (nouns formed by adding suffixes to verbs or adjectives, such as “justification”, “development”, “robustness”)
  • Agentless passive voice (such as “The model was fitted”)
  • Hedging phrases (such as “at about”, “something like”, “almost”)
  • Phrasal coordination (such as using “and” to coordinate between noun phrases)
  • Clausal coordination (such as using “and” at the beginning of clauses).
A full list of features is provided in the Appendix, Table LABEL:tbl:all-biber-features. Combined, these variables give a stylistic sense of how these texts read. These stylistic profiles, in turn, can enrich our understandings of the affordances and limitations of LLMs as they are currently developed and trained.
To tag the texts, we used the pseudobibeR package for R, which counts Biber features present in dependency-parsed texts. [^3] Available at https://github.com/browndw/pseudobibeR pseudobibeR has proven effective at reproducing results similar to those of other studies that use Biber’s tagset (DeJeu and Brown, 2023; Markey et al., 2024). Once tagged, we counted the rate of occurrence of each feature per 1,000 words of text. These rates served as the features for our analysis.
Corpus linguists studying language use and variation have regularly turned to counts of morphosyntactic and lexical features as the basis for their quantitative analysis, an approach which dominates the field (Biber, 1988; Brezina, 2018; Hyland, 2000; Wallis, 2021). Biber’s tagset offers a compelling approach to studies of linguistic variation: because linguistic features both routinely co-occur and are sensitive to context (DeJeu and Brown, 2023), analysis of feature groupings that vary across sources lends insight into the salient distinctions between the stylistic profiles of human and LLM texts. We chose this tagset to represent this perspective, but also because it has been validated by a considerable amount of scholarship (Friginal, 2013; Hardy and Römer, 2013). Of course, other foundational corpus-based studies have used different feature sets or “bag-of-words” methods, generating word frequency lists to investigate keywords, or for instance, markers of stance in academic writing (Aull and Lancaster, 2014; Hyland, 2005), Instead, Biber’s morphosyntactic tagset presents a more robust method for analyzing register and genre, capturing a wide range of linguistic functions.
2.4 Classification
To evaluate whether the features described above are sufficient to distinguish between human- and LLM-written text, we randomly split the HAP-E documents into a training set (75%) and a test set (25%), stratifying by document category.
First, we fit a random forest to classify the source of each of the texts using the Biber features. This is a seven-class classification task (human and six LLMs). This random forest identified the features most important for distinguishing between sources; these top features will be used in Section 3 for visualizations of feature variations by source. This was fit using the ranger package (Wright and Ziegler, 2017) with its default settings.
Next, we trained a random forest for each pairwise human vs. LLM training task: human vs. GPT-4o, human vs. Llama 3 8B, and so on. For comparison, we also fit lasso-penalized logistic regression models using the same set of Biber features. We used the performance of the lasso models to benchmark and provide context to the random forest models. Since we only account for linear additive contributions of the Biber features in the lasso models, any noticeable improvements in performance of the random forests are due to the flexibility of tree-based models in capturing the presence of non-linear interactions. All lasso logistic regression models were fit using the glmnet package with the one-standard-error regularization penalty selected with 10-fold cross-validation (Friedman et al., 2010).
Finally, we used these models to classify text from the M4 parallel corpus released by Wang et al. (2024). We focus on its arXiv GPT-3.5 dataset, which includes both real abstracts of preprints posted on arXiv and text generated by GPT-3.5 when prompted to write an abstract based on the preprint’s title. This provides an external test set, generated on a different task with a different model, to evaluate the generalizability of our classifiers.
Detection has been a particular emphasis since the release of ChatGPT and similar LLMs for understandable reasons (Liu et al., 2024; Waltzer et al., 2024). While our work involves similar classification tasks, we use classification not as an end in and of itself, but rather as method for helping us understand how LLMs write: how the text that they produce compares to the writing humans produce and how the text produced by one LLM compares to that produced by another.
3 Results
3.1 Classifying text by source
Figure 2: Confusion matrix for the seven-class random forest using HAP-E, evaluated on the test set. The block diagonal structure indicates that most classification errors were between different versions of the same LLM, rather than between humans and LLMs.
The random forest classifier for the seven-task problem (distinguishing between human chunk 2 and the six LLMs) achieved a test accuracy of 66%, compared to an expected accuracy of 14% from random guessing. The confusion matrix, shown in Figure 2, demonstrates that little of the error was due to confusion between human texts and the LLMs: instead, most classification errors confused Llama 3 8B and 70B, 8B Instruct and 70B Instruct, or GPT-4o and 4o Mini. Each pair consists of models of two different sizes trained on similar data, implying that the size difference does not produce dramatically different style. Overall, only 4.2% of LLM texts were falsely classified as human, and only 9.8% of human texts were falsely classified as LLMs. Feature importances for all 66 features are shown in Table LABEL:tbl:top-features in the Appendix.
3.2 Differences in style and vocabulary
Figure 3 illustrates the large variation in rate of occurrence of the fifteen most important features, relative to the rate observed in the human text; rates for all 66 features are provided in Table LABEL:tbl:top-features in the Appendix. All four instruction-tuned models have strong preferences for present participial clauses, ‘that’ clauses as subjects, nominalization, and phrasal co-ordination, which are typical markers of more informationally dense, noun-heavy style of writing. There are also signs of “house styles”: both GPT-4o models avoid clausal coordination, while all Llama 3 variants use it more frequently than humans; while both GPT-4o models use downtoners (such as barely or nearly) more frequently than humans, all Llama 3 variants avoid them.
Figure 3: Rate of Biber feature use by different LLMs, relative to the human usage of each feature, for the top 15 most important features. Note the log scale. GPT-4o and GPT-4o Mini show the largest variation from human texts, while the base variants of Llama 3 most closely resemble human grammar and style. Larger models (top row) generally show the same rhetorical differences as their smaller counterparts (bottom row), despite performing better on other benchmark tasks.
One might expect larger models to better match human text than smaller models (e.g. Llama 70B versus Llama 8B, or GPT-4o versus GPT-4o Mini), but this does not appear to be the case in Figure 3. Also, instruction tuning appears to make the model output less human, not more: the Llama 3 base models use features at rates similar to human texts, while GPT-4o and Llama 3 instruction-tuned models have much wider variation from feature to feature.
Figure 4: Rates of word use by different LLMs (per 1,000 words) compared to the human use of each word in chunk 2 (log scale). Includes all words used more than once per million words in chunk 2. Words are lemmatized to group together inflected forms. Dashed blue lines indicate the range between $10\times$10 × more and $10\times$10 × less than human use. Note that the instruction-tuned models show more variation from the diagonal, indicating more deviation in vocabulary use relative to humans.
Similar to past research (Liang et al., 2024), we find that LLMs also favor specific vocabulary. Figure 4 shows the rate of usage for words used more than once per million words by humans, comparing the usage of each LLM to the usage by humans in Chunk 2 of HAP-E. Compared to the base Llama models, in the instruction-tuned Llama and GPT-4o models certain words get used at dramatically higher and lower rates. For example, GPT-4o and 4o Mini use words like camaraderie, palpable, tapestry, and intricate at more than 100 times the rate of humans, such as in the GPT-4o output phrase “The camaraderie was palpable.” Conversely, they use certain obscenities more than 100 times less often. The top overrepresented and underrepresented words are presented in Tables LABEL:tbl:overrepresented and LABEL:tbl:underrepresented in the Appendix. [^4] Some overuse may be artifacts of the generation process; for example, Llama 3 instruction-tuned variants overuse continuation because their outputs sometimes begin with “Here is the continuation of the text…” Llama base models have a tendency to repeat themselves, so Llama 3 8B uses Deborah at 52 times the rate of humans largely because of a single document repeating it 308 times.
While many words listed in Table LABEL:tbl:overrepresented may be occasionally expected in belletristic works of fiction, their pervasiveness across LLM output in a diverse array of genres is notable. To those familiar with academic writing, newspapers, or television scripts, these words are largely unexpected, and to experts likely signal an overwritten, sentimental, or simply uneven text. In the GPT models in particular, many of these words connote some form of complex relation among objects (e.g., tapestry, intricate, camaraderie, cacophony, amidst). Coupled with positive items such as vibrant and solace, these words together may signal a preference for grandiose, if hollow, summative sentences. As the use is most similar to humans for the base models, this suggests the word choice bias is introduced by the instruction tuning process, not simply by bias in the texts composing the training sets.
3.3 Distinguishing individual LLMs
When classifying between human-generated text and one specific LLM, rather than comparing all LLMs, our classifiers achieve much higher accuracy. Typical accuracies achieved by random forests were 93–98%, as shown in Table LABEL:tbl:rf in the Appendix, even when trained on HAP-E and tested on CAP or vice versa. The lasso-penalized logistic regression classifiers attained similar performance for all LLMs except for the Llama base models, which had accuracies around 75%, as shown in Table LABEL:tbl:lasso. Since the lasso regressions only consider additive terms, this implies that interactions between the Biber features contain relevant signals for the Llama base models.
For both methods, the lower classification accuracy for the Llama base models relative to the GPT-4o and instruction-tuned Llama models indicates that instruction tuning may lead to writing that is easier to distinguish from human writing.
3.4 Generalization across corpora
When each pairwise random forest was used to classify arXiv preprints from the M4 corpus (Wang et al., 2024), accuracy dropped significantly. Models trained on instruction-tuned LLMs were able to classify M4’s GPT-3.5 output with greater-than-chance accuracy, but models trained on the Llama base outputs attained only 50% accuracy, equal to random guessing. Full results are presented in Tables LABEL:tbl:hape-v-arxiv and LABEL:tbl:coca-v-arxiv. These results demonstrate that instruction-tuned LLMs do have features in common that permit their classification, but that generalizability across LLMs or to different registers of text is difficult.
4 Discussion
This study identifies salient differences both between human and LLM-generated texts and among various models. The features that distinguish between humans and different LLMs include present participial clauses, ‘that’ clauses as sentence subject, passive voice, and nominalizations, to name a few. These findings corroborate other research that points out the ways these features produce informationally dense prose (Markey et al., 2024). Other research links these features to increased lexical diversity in generated text, as well as human judgments of linguistic mastery (Herbold et al., 2023). Lastly, Jiang and Hyland (2024) found that ChatGPT-4 text evidences more nominalizations, and fewer human subjects and epistemic stance markers, findings we see reproduced in our list of distinguishing features.
A second major finding of this research is the apparently central role of instruction tuning in creating these discrepancies between human and model general texts. While we do not have access to untuned versions of GPT, comparisons between Llama’s base and tuned models emphasize the degree to which instruction tuning pushes models to produce text that reads unlike a human. This suggests that differences in style are not simply due to the selection of texts for training the base models, but due to human preferences in instruction tuning, either deliberately or as an unintended consequence of other goals. To our knowledge, this is not discussed elsewhere in similar research.
A third major finding of this work is significant success of Biber’s tagset in modeling and classifying text. This success suggests that varied linguistic perspectives—which, perhaps, are not prioritized during the development and in-house assessments of LLMs—can reveal otherwise tacit information that distinguishes a text as machine-generated. With the linguistic perspective offered by pseudobibeR, we built a model that recognizes machine-generated text with relative ease. Our study reveals the clear value of linguistics expertise and functional conceptions of language in both LLM use and development.
That said, our intention is not to propose another way to construct LLM detectors or to police the writing of students and learners. Instead, we maintain that this type of comparative analysis is useful for identifying differences between human- and machine-generated text, zeroing in on specific teachable moments in the revision of machine-generated text.
As LLMs are increasingly put to work completing diverse writing tasks, these results suggest a notable misalignment between generated texts and the contexts in which we put them to use. This is another way of saying that LLMs do not vary their linguistic output in response to contextual factors in ways similar to humans. For those proficient in a genre—think, a therapist collating notes on a patient or, say, a college graduate writing thank-you notes to friends and family—this misalignment may not be a problem. Previous experience, in this case, guides a current sense of what a particular document should look like in order to be successful. For experts, then, LLMs appear a worthwhile productivity tool, suitable so long as they lend their expertise to further shaping the output.
For learners, though, LLMs appear problematic. Of course, using LLMs to learn more about a concept, or help generate ideas, is one thing. Using output in a text is a different matter, one that may affect a student’s learning trajectories. In this case, students offload the important cognitive labor of shaping a text for a particular audience and purpose to the LLM. Never mind that LLMs do not appear to write like humans—when students offload writing work, they offload opportunities to learn how to write IMRaD articles, client reports, executive summaries, investment pitches, etc. When LLMs are used in the classroom as writing tools, instructors of all levels and disciplines need to help students see both the shortcomings of the generated text and avenues for improvement. LLMs are not bad, either technically or morally—it is only that instructors must help inculcate in students the critical perspective of the expert to know what’s working and what isn’t.
In other contexts, however, overreliance on LLMs could produce output that might be awkward and inauthentic (e.g., in a creative genre), confusing (e.g., in instructional material), or unpersuasive (e.g., in argumentative texts). The current work thus suggests the importance of LLM practice—both in and out of the classroom—informed by human expertise via a continual dialogue of creation and revision, where LLM users are more aware and mindful of the effective uses as well as the limitations of various LLMs.
Acknowledgments
We thank members of the TeachStat Research Group for helpful discussions, the Dietrich College of Humanities and Social Sciences at Carnegie Mellon University for use of the Wright GPU cluster, and Aadi Menon for exploring suitable prompts.
Data availability
The Human-AI Parallel English (HAP-E) Corpus described in Section 2.1 is available at:
The Corpus of Contemporary American English (COCA) can be obtained at https://www.corpusdata.org/. Due to the licensing agreement, we cannot redistribute the text, but Biber features for the subset analyzed above can be found at https://huggingface.co/datasets/browndw/coca-ai-parallel-corpus-biber
Appendix A Appendix
Table 2: Lexical, grammatical, and rhetorical features detected by pseudobibeR, adapted from Biber (1988, Appendix II).
Feature Description Examples
Past tense Verbs in past tense I ran far.
Perfect aspect Verbs in perfect aspect I have written this sentence.
Present tense Verbs in present tense. the duck walks quickly
Place adverbials Adverbs and adverbial phrases describing place above, beside, outdoors
First-person pronouns I, we, our
Second-person pronouns you, your
Third-person pronouns (excluding it) he, she, they, their
Pronoun it it, its, itself
Demonstrative pronouns Pronouns replacing nouns That is an example sentence.
Indefinite pronouns anybody, nothing, someone
Pro-verb ‘do’
‘Wh-’ questions Direct who, what, when, where, and why questions When are you leaving?
Nominalizations Nouns ending in -tion, -ment, etc. justification, development, robustness
Gerunds Participial forms functioning as nouns
Other nouns All other nouns
Agentless passives Agentless passive voice The model was fitted
‘By-’ passives Passive voice with agent The task was done by Steve
‘Be’ as main verb Use of “be” forms as main verb
Existential ‘there’ “There” used to assert something exists There is a feature in this sentence
‘That’ verb complements I said that he went
‘That’ adjective complements I’m glad that you like it
‘Wh-’ clauses Clauses beginning with ‘wh-’ words (who, what, when, …) I believed what he told me
Infinitives Uninflected verb preceded by to she tried to explain
Present participial clauses Adverbial clauses used as present participles Stuffing his mouth with cookies, Joe ran out the door
Past participial clauses Adverbial clauses used as past participles Built in a single week, the house would stand for fifty years
Past participial postnominal Reduced relative past participial clauses The solution produced by this process
Present participial postnominal Reduced relative present participial clauses The event causing this decline
‘That’ clauses as subject ‘That’ relative clauses in subject position the dog that bit me
‘That’ clauses as object ‘That’ relative clauses in object position the dog that I saw
‘Wh-’ relatives as subject ‘Wh-’ relatives in subject position the man who likes popcorn
‘Wh-’ relatives as object ‘Wh-’ relatives in object position the man who Sally likes
Pied-piping relative clauses Relative clauses moved in sentence by ‘wh-’ questions the manner in which he was told
Sentence relatives Bob likes fried mangoes, which is the most disgusting thing I’ve ever heard of
Because Causative adverbial subordinator because
Though Concessive adverbial subordinators although, though
If, unless Conditional adverbial subordinators if, unless
Other adverbial subordinators since, while, whereas
Prepositional phrases
Attributive adjectives the big horse
Predicative adjectives the horse is big
Adverbs Total adverbs
Type-token ratio Type-token ratio, including punctuation
Mean word length Average word length in characters, excluding punctuation
Conjuncts consequently, furthermore, however
Downtoners barely, nearly, slightly
Hedges at about, something like, almost
Amplifiers absolutely, extremely, perfectly
Emphatics a lot, for sure, really
Discourse particles sentence-initial well, now, anyway
Demonstratives “That”, “this”, “these”, or “those” used as determiners That is the feature
Possibility modals can, may, might, could
Necessity modals ought, should, must
Predictive modals will, would, shall
Public verbs Verbs indicating speaking or announcing assert, declare, mention, predict, swear
Private verbs Verbs expressing a personal intellectual state assume, believe, doubt, presume, understand
Suasive verbs Verbs indicating persuasion command, insist, propose
‘Seem’ and ‘appear’ Used as verbs
Contractions can’t, won’t
‘That’ deletion ‘That’ omitted as subordinator I think [that] he went
Stranded prepositions Preposition at the end of sentence or clause rather than before its object she is the candidate that I was thinking of
Split infinitives Infinitive with adverb between to and the verb he wants to convincingly prove that…
Split auxiliaries they were apparently shown to…
Phrasal co-ordination Pairs of nouns, verbs, adjectives, or adverbs connected by a coordinating conjunction The nouns and verbs are coordinated
Clausal co-ordination Sentence clauses coordinated by a conjunction The sentence was long, but I read it anyway
Synthetic negation no answer is good enough
Analytic negation that isn’t good enough
Table 3: Biber features distinguishing human- and LLM-written text in HAP-E Rate per 1,000 tokens; LLM rates relative to human Chunk 2. Blue indicates higher-than-human usage, red indicates lower. Ordered by feature importance for the seven-class classification task.
Human GPT Llama 3 Instruct Llama 3 Base
Feature Chunk 1 Chunk 2 GPT-4o Mini GPT-4o 8B 70B 8B 70B Importance
Present participial clauses 1.7 1.7 481% 527% 224% 261% 94% 102% 1,715.1
‘That’ clauses as subject 2.1 2.1 331% 263% 180% 173% 64% 68% 1,387.2
Mean word length 4.5 4.4 114% 116% 101% 103% 99% 100% 1,070.4
Adverbs 67.9 71.8 86% 82% 73% 75% 102% 102% 916.0
Nominalizations 14.6 14.6 209% 214% 145% 151% 88% 91% 808.9
Phrasal co-ordination 6.7 6.1 144% 194% 187% 170% 92% 97% 776.0
Clausal co-ordination 11.2 12.4 63% 59% 141% 127% 120% 116% 727.3
Other nouns 254.9 240.6 97% 103% 91% 95% 91% 94% 725.5
Downtoners 1.9 1.9 155% 118% 60% 57% 68% 73% 712.8
Attributive adjectives 46.7 43.8 140% 150% 100% 104% 79% 83% 677.7
Demonstratives 6.2 6.5 137% 133% 77% 80% 75% 80% 675.9
Agentless passives 7.7 7.8 51% 53% 96% 89% 101% 98% 664.4
Past participial postnominal 1.6 1.5 257% 235% 75% 75% 129% 131% 633.7
Infinitives 15.9 16.5 87% 83% 140% 132% 120% 113% 614.9
Prepositional phrases 101.0 98.0 118% 118% 100% 100% 87% 90% 581.2
Emphatics 8.6 9.2 75% 68% 97% 98% 76% 75% 545.9
Existential ‘there’ 1.9 2.1 59% 71% 42% 42% 108% 109% 527.3
Possibility modals 5.3 5.7 144% 104% 111% 116% 99% 99% 521.1
Past tense 37.5 41.9 77% 83% 91% 83% 115% 111% 519.0
‘That’ verb complements 2.1 2.5 70% 55% 144% 114% 159% 147% 503.4
Contractions 16.9 18.1 63% 60% 141% 139% 142% 129% 497.7
Other adverbial subordinators 5.9 6.2 136% 118% 114% 107% 82% 89% 496.6
‘Be’ as main verb 29.0 30.0 61% 63% 107% 100% 108% 101% 494.3
Public verbs 5.8 6.8 53% 63% 65% 67% 119% 112% 466.1
Present participial postnominal 1.4 1.3 293% 243% 125% 124% 101% 113% 451.0
Predicative adjectives 5.8 6.2 84% 90% 140% 131% 169% 154% 450.8
Suasive verbs 2.9 3.2 98% 116% 89% 90% 141% 137% 449.6
Analytic negation 8.5 9.7 73% 61% 80% 78% 113% 107% 439.6
Third-person pronouns 26.2 29.8 91% 91% 108% 104% 123% 120% 436.8
Private verbs 16.0 18.1 91% 85% 122% 113% 128% 126% 435.9
Perfect aspect 7.5 7.2 62% 60% 121% 111% 96% 92% 434.2
Pro-verb ‘do’ 2.9 3.2 25% 26% 60% 59% 115% 109% 430.8
Present tense 57.2 58.3 82% 76% 99% 100% 119% 116% 418.9
Amplifiers 2.0 2.1 63% 46% 40% 45% 85% 85% 412.6
Pronoun ‘it’ 12.5 13.2 88% 92% 105% 105% 108% 105% 412.0
Gerunds 3.0 3.0 156% 152% 119% 124% 83% 86% 402.9
First-person pronouns 34.0 35.3 81% 62% 111% 108% 136% 127% 393.4
Because 1.2 1.5 19% 20% 38% 38% 121% 105% 387.9
Split auxiliaries 3.4 3.4 91% 77% 126% 118% 98% 92% 379.4
Time adverbials 3.7 3.7 77% 79% 68% 70% 115% 119% 366.0
Though, although 0.5 0.5 65% 129% 21% 28% 83% 90% 361.2
Place adverbials 3.4 3.4 146% 144% 101% 99% 118% 128% 353.9
Demonstrative pronouns 5.5 6.1 55% 50% 71% 76% 99% 97% 350.8
Predictive modals 5.1 5.6 72% 57% 95% 106% 118% 111% 345.4
Indefinite pronouns 3.2 3.3 73% 77% 84% 83% 156% 149% 323.5
by-passives 0.9 0.9 56% 63% 118% 107% 113% 103% 314.9
Synthetic negation 1.2 1.3 36% 51% 36% 36% 93% 91% 311.8
‘Wh-’ questions 1.3 1.4 89% 56% 141% 137% 140% 132% 306.1
If, unless 2.2 2.6 78% 60% 58% 62% 117% 111% 305.1
‘Seem’ and ‘appear’ 0.7 0.7 131% 179% 140% 128% 99% 105% 290.3
Necessity modals 1.1 1.3 105% 78% 51% 54% 143% 137% 286.7
Conjuncts 2.1 2.3 110% 96% 97% 116% 99% 104% 281.2
Second-person pronouns 14.9 15.7 63% 52% 77% 81% 118% 110% 277.3
Stranded prepositions 0.8 0.9 66% 66% 111% 114% 87% 81% 266.5
‘Wh-’ relatives as subject 2.3 2.1 70% 66% 109% 102% 106% 103% 259.1
Sentence relatives 1.0 1.0 50% 51% 104% 117% 101% 92% 251.7
‘Wh-’ clauses 1.7 1.9 75% 66% 90% 78% 126% 122% 247.1
Discourse particles 1.1 1.0 60% 60% 89% 91% 146% 140% 181.9
Hedges 1.1 1.3 50% 63% 62% 67% 89% 92% 181.1
Past participial clauses 0.3 0.3 273% 307% 158% 150% 59% 71% 171.7
‘That’ deletion 0.7 0.8 75% 66% 75% 73% 134% 151% 169.8
Pied-piping relative clauses 0.6 0.6 59% 56% 38% 43% 50% 54% 164.5
‘That’ clauses as object 0.6 0.6 65% 56% 66% 60% 93% 90% 118.3
‘That’ adjective complements 0.3 0.4 80% 51% 121% 106% 148% 139% 107.6
Split infinitives 0.2 0.3 79% 94% 110% 104% 44% 49% 94.9
‘Wh-’ relatives as object 0.3 0.3 20% 13% 8% 13% 61% 62% 62.5
Table 4: Most overrepresented words in LLM texts Rates relative to human chunk 2
Llama 3 8B Llama 3 8B Instruct Llama 3 70B Llama 3 70B Instruct GPT-4o Mini GPT-4o
Word Rate Word Rate Word Rate Word Rate Word Rate Word Rate
deborah 52 unease 101 bananas 31 unease 63 camaraderie 171 camaraderie 162
rambo 22 continuation 52 paperback 30 palpable 47 tapestry 147 tapestry 155
matty 20 palpable 48 bam 26 continuation 29 palpable 145 intricate 119
goodnight 18 reminder 33 verona 25 shoutout 28 grapple 131 underscore 107
ml 15 pang 29 filth 19 intricate 27 intricate 129 unspoken 102
merlin 13 rut 29 rekall 17 pang 25 fleeting 124 amidst 100
worcester 11 waft 28 denis 14 camaraderie 24 ignite 122 palpable 95
fay 10 prioritize 27 darry 12 policymaker 24 vibrant 92 solace 95
missy 10 grapple 24 ebook 12 prioritize 24 amidst 90 fleeting 84
elisa 10 camaraderie 23 janice 12 reminder 24 cacophony 89 unravel 83
Table 5: Most underrepresented words in LLM texts Rates relative to human chunk 2
Llama 3 8B Llama 3 8B Instruct Llama 3 70B Llama 3 70B Instruct GPT-4o Mini GPT-4o
Word Rate Word Rate Word Rate Word Rate Word Rate Word Rate
bingo 0.033 anyways 0.0097 jeez 0.040 yep 0.0075 extremely 0.0034 i.e. 0.0026
scorch 0.040 analyse 0.011 donnie 0.042 fucking 0.0075 spokesman 0.0059 blah 0.0063
been 0.043 somebody 0.012 alexithymia 0.046 horrible 0.0077 i.e. 0.0095 fuck 0.0079
abt 0.043 characterise 0.013 frampton 0.048 ok 0.011 unhappy 0.010 fucking 0.0083
bananas 0.043 fuck 0.014 y’all 0.048 fuckin 0.013 bitch 0.012 asshole 0.0093
pham 0.043 obviously 0.014 analogous 0.053 amid 0.015 fucking 0.012 nasty 0.011
ie 0.045 ok 0.014 carotenoid 0.053 fortunately 0.017 kg 0.013 chum 0.013
monstrous 0.045 fuckin 0.014 mussel 0.053 i.e. 0.018 fuck 0.014 shit 0.014
pow 0.045 blah 0.015 que 0.053 obviously 0.018 visa 0.015 ok 0.014
unquote 0.045 yep 0.016 whatsapp 0.053 amongst 0.019 ter 0.016 and/or 0.016
Table 6: Word counts in the HAP-E corpus, by source
Author acad (n = 1227) blog (n = 1526) fic (n = 1395) news (n = 1322) spok (n = 1721) tvm (n = 1099) Total
Human
       Chunk 1 602,878 748,901 674,437 644,896 833,699 598,157 4,102,968
       Chunk 2 602,683 749,052 675,595 646,030 832,797 591,704 4,097,861
ChatGPT
       GPT-4o 633,884 826,590 759,461 708,375 969,210 606,059 4,503,579
       GPT-4o Mini 690,828 916,433 837,527 771,324 1,036,125 678,429 4,930,666
Llama Base
       Llama 3 70B 510,942 731,889 651,557 572,423 1,098,698 502,012 4,067,521
       Llama 3 8B 544,845 735,068 740,610 510,671 1,083,953 525,037 4,140,184
Llama Instruction-Tuned
       Llama 3 70B Instruct 580,579 735,501 641,982 627,299 934,911 453,559 3,973,831
       Llama 3 8B Instruct 5550,463 679,884 599,000 568,703 836,531 448,874 3,683,455
Total 4,717,102 6,123,318 5,580,169 5,049,721 7,625,924 4,403,831 33,500,065
Table 7: Word counts in the CAP corpus, by source
Author acad (n = 1221) blog (n = 1071) fic (n = 1372) mag (n = 1293) news (n = 1188) spok (n = 1334) tvm (n = 1006) web (n = 1130) Total
Human
       Chunk 1 594,011 521,050 669,470 629,726 579,755 651,956 497,733 549,341 4,693,042
       Chunk 2 593,875 521,493 671,400 630,126 580,208 651,013 497,900 549,736 4,695,751
ChatGPT
       GPT-4o 646,061 578,298 754,715 696,988 640,690 734,320 562,464 609,670 5,223,206
       GPT-4o Mini 710,560 631,028 830,929 759,785 697,919 793,067 608,491 664,693 5,696,472
Llama Base
       Llama 3 70B 634,881 517,613 739,502 681,171 583,279 747,996 464,058 570,922 4,939,422
       Llama 3 8B 625,065 527,628 729,348 686,137 553,703 736,093 462,948 550,569 4,871,491
Llama Instruction-Tuned
       Llama 3 70B Instruct 607,677 525,063 637,822 632,915 561,171 694,292 406,103 549,999 4,615,042
       Llama 3 8B Instruct 554,642 480,391 597,560 572,409 514,765 624,254 409,802 508,688 4,262,511
Total 4,966,772 4,302,564 5,630,746 5,289,257 4,711,490 5,632,991 3,909,499 4,553,618 38,996,937
Table 8: Random forest pairwise classification accuracy, distinguishing each LLM from human text. Training accuracy based on out-of-bag error
Trained on HAP-E Trained on CAP
LLM Training Test on CAP Training Test on HAP-E
GPT-4o
       GPT-4o 98.3% 98.4% 98.9% 97.4%
       GPT-4o Mini 98.7% 98.2% 99.2% 97.5%
Llama Instruct
       Llama 3 70B Instruct 94.5% 93.5% 94.6% 91.8%
       Llama 3 8B Instruct 95.5% 95.3% 96.1% 94.1%
Llama Base
       Llama 3 70B 93.0% 94.6% 95.8% 89.1%
       Llama 3 8B 94.0% 93.6% 95.6% 89.4%
Table 9: Lasso pairwise classification accuracy, distinguishing each LLM from human text
Trained on CAP Trained on HAP-E
LLM Training Test on HAP-E Training Test on CAP
GPT-4o
       GPT-4o 95.9% 95.3% 96.5% 95.3%
       GPT-4o Mini 96.7% 96.3% 96.9% 96.2%
Llama Instruct
       Llama 3 70B Instruct 91.1% 90.7% 92.1% 90.2%
       Llama 3 8B Instruct 92.4% 91.8% 92.9% 91.9%
Llama Base
       Llama 3 70B 77.2% 73.2% 75.6% 75.3%
       Llama 3 8B 77.6% 73.8% 76.3% 75.3%
Table 10: Random forest accuracy when trained on HAP-E and classifying M4 arXiv data.
Model Acc. (in-sample) Acc. (out-of-sample)
ChatGPT
       GPT 4o 98.22% 70.63%
       GPT 4o Mini 98.73% 60.77%
Llama Instruction-Tuned
       Llama 3 70B Instruct 95.33% 68.20%
       Llama 3 8B Instruct 96.20% 65.60%
Llama Base
       Llama 3 70B 92.43% 51.17%
       Llama 3 8B 93.56% 51.85%
Table 11: Random forest accuracy when trained on CAP and classifying M4 arXiv data.
Model Acc. (in-sample) Acc. (out-of-sample)
ChatGPT
       GPT 4o 98.76% 70.68%
       GPT 4o Mini 99.04% 65.27%
Llama Instruction-Tuned
       Llama 3 70B Instruct 95.35% 59.27%
       Llama 3 8B Instruct 96.60% 57.95%
Llama Base
       Llama 3 70B 95.53% 50.92%
       Llama 3 8B 95.39% 51.48%